OpenStudy (anonymous):

who illustrated x+(-x=0

OpenStudy (anonymous):

x+(-x)=0 x - x =0 any number minus itself is zero

OpenStudy (skullpatrol):

logic illustrates it

OpenStudy (anonymous):

that is "what"

OpenStudy (anonymous):

I didn't know that this particular thing has an author...

OpenStudy (skullpatrol):

It has no "author", it is a given property of all real-numbers.

OpenStudy (anonymous):

oops clicked something, don't know what. I was just reading from the user's question. Based on it, x+(-x) has some sort of an author.

OpenStudy (anonymous):

well, doesn't matter so much to me. bye

OpenStudy (skullpatrol):

Property of opposites: a + (-a) = 0.

OpenStudy (skullpatrol):

we need it to talk about negative numbers :-)

OpenStudy (skullpatrol):

a + (-a) = 0 is a direct logical consequence of the reflexive property of equality: a = a.

OpenStudy (anonymous):